Assuming the mean and SD are the same for this year’s national exam, what percentage of test-takers scored above and below you if your score is 1900 and given that X ~ N(1600, 140). Using the empirical rule, find the percentage of students who scored between 1880 and 1920?


1
Expert's answer
2021-12-16T07:32:46-0500

Given that "X\sim N(1600, 140)."

Then



"\mu=1600, \sigma^2=140, \sigma=\sqrt{140}""P(X<1900)=P(Z<\dfrac{1900-\mu}{\sigma})""=P(Z<\dfrac{1900-1600}{\sqrt{140}})\approx P(Z<25.3546)""\approx1(100\%)"

The empirical rule formula:

68% of data falls within 1 standard deviation from the mean - that means between "\u03bc - \u03c3"  and "\u03bc + \u03c3."

95% of data falls within 2 standard deviations from the mean - between "\u03bc \u2013 2\u03c3"  and "\u03bc + 2\u03c3."

99.7% of data falls within 3 standard deviations from the mean - between "\u03bc - 3\u03c3" and "\u03bc + 3\u03c3."

Given



"\mu=1600, \sigma=\sqrt{140}\approx11.83"

Then



"(\mu-\sigma, \mu+\sigma)\approx(1588.2, 1611.8)""(\mu-2\sigma, \mu+2\sigma)\approx(1576.3, 1623.7)""(\mu-3\sigma, \mu+3\sigma)\approx(1564.5, 1635.5)"

We cannot use the empirical rule to find the percentage of students who scored between 1880 and 1920.